OpenStudy (anonymous):

what is -5/6 divided by 9/10? A. -27/25 B. -4/3 C. -3/4 D. -25/27

OpenStudy (anonymous):

D :)

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (mysangels):

\[-5/6 \div 9/10 = -5/6 \times 10/9\]

OpenStudy (mysangels):

Then you would multiply, and end up with -50/54 Now you would simplify

OpenStudy (anonymous):

thank you too! I will fan you both.